Lactic acid, HC3H5O3, is a substance found in sour milk products such as yogurt, and is...

  1. Lactic acid, HC3H5O3, is a substance found in sour milk products such as yogurt, and is produced naturally by fermentation in body cells during normal metabolism and exercise The Ka for lactic acid is 38 x 10-4. Its sodium salt, sodium lactate, NaC3H5O3, is used in foods as a preservative and acidity regulator. It is also used in shampoo products as an effective humectant and moisturizer. Determine the pH of a solution prepared by adding 4.5 grams of sodium lactate and 2.0 grams of lactic acid to 500 ml of total solution.
Answer #1

Total volume of buffer solution = 500mL =0.5L

mass of acid = 2.0g

molar mass of acid = 90g/mol

molarity of acid = moles/V(L) = [mass/molar mass] /V(L)

= [2.0g/90g/mol] /0.5 L =0.044 M

mass of salt = 4.5g

molar mass of salt = 112 g/mol

molarity of salt = [4.5g/112g/mol] / 0.5L = 0.0803 M

Ka of acid = 3.8x10-4

pKa of acid = -log 3.8x10-4 = 3.42

The pHof an acidic buffer is given by Hendersen equation as

pH = pKa + log [conjugate base/salt] / [acid]

= 3.42 + log (0.0803/0.044)

= 3.682